# Let Guidelines Guide You: A Prescriptive Guideline-Centered Data Annotation Methodology

### v1

@@ 1 Introduction
In recent years, several works have highlighted the importance of clear guidelines for producing high-quality data annotations. In particular, researchers have framed distinct annotation processes into annotation paradigms, each with its own set of characteristics, limitations, and requirements. The most notable contributions in this field are the prescriptive and descriptive annotation paradigms, originally designed to address subjective tasks in NLP Rottger et al. (2022), such as hate speech detection, subjectivity detection, and sentiment analysis. The prescriptive annotation paradigm enforces a shared belief of the task conveyed into a unique set of annotation guidelines on which annotators are instructed. In contrast, the descriptive annotation paradigm favors individual perspectives and beliefs of the annotators, which are allowed to be reflected in the task. Another notable emergent annotation paradigm is the perspectivist one Cabitza et al. (2023), where multiple annotation ground-truths for the same data sample are allowed while sharing a unique set of annotation guidelines. All these paradigms share the underlying motivation of addressing annotation issues, leading to low-quality data. These issues are largely evident in subjective tasks like hate speech detection Plank et al. (2014); Sap et al. (2022), and subjectivity detection Riloff and Wiebe (2003); Savinova and Moscoso Del Prado (2023), among others.
@@
In a standard annotation process according to the prescriptive paradigm, annotators are tasked to annotate data samples to a class set given a set of guideline definitions. The main issue we identify in this process is that annotators interpret and apply annotation guidelines implicitly, i.e., without reporting directly the mapping between classes and guidelines. This introduces a disconnection between annotated data and the intended guidance encapsulated within the guidelines, thus allowing annotators to align a data sample to the class set by leveraging their own beliefs. Hereinafter, we refer to this concept as annotation shortcuts. The adherence of the annotation process to annotation guidelines cannot be analyzed via the reported class sets where apparent non-conflicting situations hide more nuanced dynamics. For instance, two annotators who agree on labeling a data sample with a particular class might reach the same conclusion by leveraging different guideline definitions. This is in contrast to the main requirement of the prescriptive paradigm, which states that annotators must adhere to the same set of guidelines. Similarly, implicit adherence to annotation guidelines limits crucial annotation phases, such as annotators’ discussion and edge case resolution Braun (2023). In particular, the risk of annotation ’shortcuts’ and annotator disagreement with respect to the guideline set also pertain to these phases.
Lastly, we broaden the impact of implicitly adhering to annotation guidelines to model training. In particular, one of the main underlying intents of training a model with annotated data is to mimic human behavior in aligning data samples to a class set Hastie et al. (2009). However, while the human annotation process is grounded on a guideline set depicting domain-specific knowledge about a task, this information is not provided to the model during training. This denotes an inconsistency between how annotators are expected to perform data annotation by relying on a set of guideline definitions and how a model is trained solely relying on the reported class.
We propose a shift in the annotation process, where annotators explicitly leverage the definitions delineated in validated annotation guidelines rather than relying on less informative class sets. In particular, we introduce a novel annotation methodology, denoted Guideline-Centered (GC), centered on designing an annotation process that adheres transparently to employed annotation guidelines. GC opens a wide spectrum of advantages that affect the way data is annotated and eventually used for model training.
By aligning the annotation process with well-defined annotation guidelines, we ensure that annotators explicitly utilize the guidance provided by annotation guidelines. This design choice limits annotation shortcuts derived from individual beliefs, resulting in a more objective and standardized annotation process. Furthermore, by aligning guideline definitions to annotated samples, GC allows training models in a way that is coherent with the annotation process carried by humans and where textual annotation guideline definitions can be used as a richer form of guidance Wang et al. (2018); Pappas and Henderson (2019); Wang et al. (2023). Similarly, the informative content provided by the annotation guidelines can benefit model evaluation, leading to a more comprehensive assessment of model performance.
Lastly, an implied consequence of GC is the definition of an annotation process that is agnostic of the employed class set. More precisely, annotations aligned with a guideline set can be mapped to any class set that relies on that guideline set. This lifts annotators from the cost of re-annotating collected data when shifting to a different class set and allows decoupling data annotation from conflict resolution on the class set, a common situation in subjective tasks Caselli et al. (2020); Antici et al. (2024).
The remainder of this paper is structured as follows. In Section 4, we frame our contribution with respect to the existing literature. Then, we discuss the standard annotation process under the prescriptive paradigm in Section 2, and we introduce our methodology in Section 3. Finally, we discuss some major implications of GC in Section 5, and report the concluding remarks in Section 6.
@@ 3.2 Properties
Figure 2: GC annotation process allows defining different class sets ($\mathcal{C}^{\alpha}$, $\mathcal{C}^{\beta}$, $\mathcal{C}^{\gamma}$) and different class grounding functions ($r^{\alpha}$, $r^{\beta}$, $r^{\gamma}$) for the same guideline set $\mathcal{G}$ at the cost of a single human annotation stage (blue box).
@@ 4 Related Work
Human evaluation practices and annotation error detection are generally understudied Krishna et al. (2023); Klie et al. (2023); Sandri et al. (2023). However, several aspects must be considered when annotating data, from best practices for designing guidelines to annotation evaluation aimed at resolving annotators’ disagreement. We overview them in detail.
Modeling the Annotation Process.
Annotation guidelines are designed based on the intended use of the collected dataset for a particular downstream task. On this basis, the annotation process can be formulated to encode one specific view or capture different beliefs about the task Rottger et al. (2022), as is often the case with subjective tasks. Different beliefs are also known as Human Label Variation Plank (2022), where systematic disagreement between annotators can be due to different factors, such as ethnicity Sap et al. (2019), sociodemographic background Sap et al. (2022); Frenda et al. (2023), or a combination of those Orlikowski et al. (2023); Garg et al. (2023). For instance, Garg et al. (2023) attempt to dissect biases for subjective tasks, specifically hate speech detection, by providing a taxonomy that could make disagreement more interpretive. However, disagreement, does not necessarily indicate the existence of an error. In an attempt to distinguish disagreement from errors, Weber and Plank (2023) proposes a human-in-the-loop Annotation Error Detection method that aims at detecting errors by repeatedly querying a human for error corrections in its prediction loop. Recognising the value of disagreement, recent research proposes to model individual annotator behavior Mostafazadeh Davani et al. (2022). For example, Sandri et al. (2023) propose a taxonomy of possible reasons leading to annotators’ disagreement in subjective tasks, suggesting that the type of disagreement should be taken into account when adjusting guidelines for a given task. Fleisig et al. (2023) take disagreement a step further and present a model that predicts individual annotator ratings on offensive text while they show that using demographic information and opinions on online content can predict annotator ratings, with non-invasive questions about online experiences reducing the need for detailed demographic data.
Reporting Annotation Guidelines.
Recent studies have explored the effectiveness of incorporating human-in-the-loop approaches to guideline design van der Stappen and Funk (2021); Antici et al. (2024); Guzmán-Monteza (2023). By involving annotators in the iterative refinement of guidelines, different views about data might be discussed and converged. For instance, when drafting guidelines for argument mining, Guzmán-Monteza (2023) identify exceptional cases that are unclear, ambiguous, or not considered, and all changes are included in the guidelines until exceptional cases do not occur anymore. This process might be facilitated by mapping the annotator’s decision to a specific guideline, as done by Oouchida et al. (2009) by proposing a framework that integrates the management of the annotation guidelines in an existing annotation tool and uses guidelines as an aid when there is disagreement. According to Lee et al. (2023), although the iterative refinement of the guidelines might increase inter-annotator agreement, it can lead to the creation of tailored annotation guidelines that can inhibit reproducibility by causing divergent annotations across different research groups. As stated by the authors, this can be mitigated by keeping track of how instances map to a specific guideline. A step towards this direction is the work of Jikeli et al. (2023), where they collect and annotate a corpus for antisemitism. In the process, they require annotators to specify which part of the definition of antisemitism applies to each instance. The authors’ goal is to reduce false positives in the automated detection of antisemitic tweets. Although focusing on a specific issue within social media content analysis, this work exemplifies the importance of keeping track of guidelines during the annotation process.

### v2

@@ 1 Introduction
In recent years, several works have highlighted the importance of clear guidelines for producing high-quality data annotations. In particular, researchers have framed distinct annotation processes into annotation paradigms, each with its own set of characteristics, limitations, and requirements. The most notable contributions in this field are the prescriptive and descriptive annotation paradigms, originally designed to address subjective tasks in NLP Rottger et al. (2022), such as hate speech detection, subjectivity detection, and sentiment analysis. The prescriptive annotation paradigm enforces a shared belief of the task conveyed into a unique set of annotation guidelines on which annotators are instructed. In contrast, the descriptive annotation paradigm favors annotators’ perspectives and beliefs, which are allowed to be reflected in the task. Another notable emergent annotation paradigm is the perspectivist one Cabitza et al. (2023), where multiple annotation ground truths for the same data sample are allowed while sharing a unique set of annotation guidelines. All these paradigms share the underlying motivation of addressing annotation issues, which lead to low-quality data and are largely evident in subjective tasks like hate speech detection Plank et al. (2014); Sap et al. (2022), and subjectivity detection Riloff and Wiebe (2003); Savinova and Moscoso Del Prado (2023).
@@
We propose a shift in the annotation process, where annotators explicitly leverage the definitions delineated in validated annotation guidelines rather than relying on less informative class sets. In particular, we introduce a novel annotation methodology, denoted Guideline-Centered (GC), focused on designing an annotation process that adheres transparently to employed annotation guidelines. GC opens a wide spectrum of advantages that affect the way data is annotated and eventually used for model training.
The remainder of this paper is structured as follows. In Section 2, we discuss the standard annotation process under the prescriptive paradigm. Then, we introduce our methodology in Section 3. In Section 4 we frame our contribution with respect to the existing literature. Finally, we discuss some major implications of GC in Section 5, and report the concluding remarks in Section 6.
@@ 3.2 Properties
Figure 2: GC annotation process allows defining different class sets ($\mathcal{C}^{\alpha}$, $\mathcal{C}^{\beta}$, $\mathcal{C}^{\gamma}$) and different class grounding functions ($r^{\alpha}$, $r^{\beta}$, $r^{\gamma}$) for the same guideline set $\mathcal{G}$ at the cost of a single human annotation stage (blue box). The annotator is unaware of the mapping between $\mathcal{G}$ and the class sets $\mathcal{C}^{\alpha}$, $\mathcal{C}^{\beta}$ and $\mathcal{C}^{\gamma}$.
@@ 4 Related Work
Annotation paradigms have been recently proposed for data annotation to encode one shared view, such as the prescriptive paradigm Rottger et al. (2022), or different beliefs about a phenomenon, such as the descriptive Rottger et al. (2022), perspectivist Cabitza et al. (2023) paradigms. The ultimate goal of annotation paradigms is to regulate data annotation for a specific intended use of collected data. In the prescriptive paradigm, this objective is achieved by means of annotation guidelines Benikova et al. (2014); Sabou et al. (2014); Zaghouani et al. (2014); Krishna et al. (2023). Notable examples address subjective tasks, such as hate speech Salminen et al. (2019); Sap et al. (2022); Davani et al. (2023), stance detection Luo et al. (2020); AlDayel and Magdy (2021), sentiment analysis Waseem and Hovy (2016); Kenyon-Dean et al. (2018), and subjectivity detection Antici et al. (2024).
In all these approaches, annotators do not keep track of the annotation guidelines they rely on to annotate data samples. Therefore, although of crucial importance, the mapping between guidelines and data samples is lost during the process. To the best of our knowledge, Jikeli et al. (2023) is the only work that designs a data annotation process that enforces explicit grounding to annotation guidelines. In particular, the authors collect and annotate a corpus for the automated detection of antisemitic tweets. During the annotation process, annotators are required to specify which part of the definition of antisemitism applies to each data sample. Our work differs from the one of Jikeli et al. (2023) in several aspects. First, the work of Jikeli et al. (2023) is limited to a specific issue within social media content analysis and does not present a general formulation of the annotation process. Second, the proposed data annotation process relies on reporting both the guideline subset $\mathcal{G}_{x}$ and class subset $\mathcal{C}_{x}$ for each data sample $x$. This process suffers from an annotation shortcut where annotators may first define $\mathcal{C}_{x}$ and subsequently $\mathcal{G}_{x}$, thus reversing the intended annotation order. Furthermore, the annotation process is tight to a specific class set $\mathcal{C}$ and has increased annotation workload due to reporting both $\mathcal{G}_{x}$ and $\mathcal{C}_{x}$. We elaborate more on these two aspects in Section 5.